OpenStudy (anonymous):

find a polynomial P(x) of degree 3 whose zeros are -2, -1, and 5.

OpenStudy (anonymous):

P(x)=(x+2)(x+1)(x-5)

hartnn (hartnn):

a polynomial with one zero as 'a' can be written as f(x)=(x-a) now, we have three zeros here, a polynomial with 3 zeros 'a','b','c' will be f(x)=(x-a)(x-b)(x-c) here, a=-2,b=-1,c=5 just plug in numbers!